OpenStudy (anonymous):

How to find the angle if 0= 3cosx - 4sinx??

OpenStudy (cwrw238):

use complementary angle method

OpenStudy (anonymous):

I kind of forgat how to use it

OpenStudy (cwrw238):

i'm stretching my memory as well....

OpenStudy (cwrw238):

i think you can use Rcos (x + a) = Rcosxcosa - Rsinxsina now compare this with 3cosx - 4sinx = 0 and compare coefficients hmm... maybe there an easier way though.. maybe you can use sinx = square root(1 - cos^2 x)

OpenStudy (cwrw238):

4 square root(1 - cos^2x) = 3 cos x 16(1 - cos^2 x) = 9cos^2 x solve for x

OpenStudy (cwrw238):

that gives cos^2x = 16/25 cos x = 4/5 or -4/5 right?

OpenStudy (anonymous):

Yeah right my textbook says the same answer,but one question how did you get the 16

OpenStudy (cwrw238):

i squared ;4 square root(1 - cos^2x)'

OpenStudy (cwrw238):

squaring both sides

OpenStudy (anonymous):

Oh okay now it makes sense to me thanks for your help

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

thats ne finished tonight - i must drink some cwrw

OpenStudy (cwrw238):

later saria

OpenStudy (anonymous):

Yeah right ....later

OpenStudy (anonymous):